Find the equation of circle passing through (4,5) and having centre at (2,2) is?

Grade:11

1 Answers

hemambarnaidu
6 years ago
centre(h,k) and radius r then eq.of circle is
(x-h)2+(y-k)2=r2
(x-2)2+(y-2)2=r2
this is passing through (4,5)
That is (4-2)2+(5-2)2=r2
4+9=13= r2
Equation of circle is (x-2)2+(y-2)2=13